What is the EVPI?

Success

Moderate Success

Failure

Probability

0.3

0.3

0.4

Sell Company

107

107

107

Form Joint Venture

210

120

89

Sell Software on own

420

234

-100

Number: Round to: 1 decimal places.
Answer #1

E(sell company) =0.3*107+0.3*107+0.4*107 =107

E(from Joint Venture )=134.6

E(sell software) =156.2

tehrefore expected value without information ,max(107,134.6,156.2) =156.2

expected value with perfect information=0.3*420+0.3*234+0.4*107=239

hence EVPI =239-156.2 =82.8